OpenStudy (anonymous):

9. Choose the equation of the vertical line passing through the point (–4, 2). (5 points) y = –4 x = –4 y = 2 x = 2 10. Choose the equation of the horizontal line that passes through the point (–5, 1). (5 points) y = –5 x = 1 y = 1 x = –5 13. Choose the equation below that represents the line passing through the point (–2, –3) with a slope of –6. (5 points) y + 3 = –6(x + 2) y + 2 = 6(x + 3) y – 3 = –6(x – 2) y – 2 = 6(x – 3)

OpenStudy (anonymous):

LETTER ANSWERS PLEASE ! 14. Choose the equation below that represents the line passing through the point (–3, –1) with a slope of 4. (5 points) y = 4x – 11 y = 4x + 11 y = 4x + 7 y = 4x – 7 15. Choose the equation of the line passing through the point (3, –2) and parallel to y = 3x – 6. (5 points) y = 3x + 7 y = 3x – 11 y = –3x + 11 y = –3x – 12 16. Choose the equation of the line passing through the point (–3, 1) and parallel to y = –x – 4. (5 points) y = x + 2 y = –x – 2 y = –x + 4 y = x – 4 17. Choose the equation of the line passing through the point (–2, 4) and perpendicular to y = –x + 4. (5 points) y = 2x + 8 y = 2x y = –2x – 8 y = –2x 18. Choose the equation of the line passing through the point (–3, –1) and perpendicular to y = x + 3. (5 points) y = –2x – 7 y = –2x + 5 y = 2x + 5 y = 2x – 7

OpenStudy (anonymous):

Wow, really, can't you do your own test?

OpenStudy (anonymous):

Well when you post your entire test you are not asking for help to understand, you are asking for help cheating

OpenStudy (anonymous):

So you are claiming that posting problems on here for other people to answer and asking only for the letter answers is not cheating? Really?

OpenStudy (anonymous):

Well you have 8 questions posted, which one do you want to start with?

OpenStudy (anonymous):

No, because I am not going to give you the answer, and you do not want to think, you only want to cheat.

OpenStudy (anonymous):

So you think that you are going to get someone else to just give you the answers.

OpenStudy (anonymous):

It's a free country, I can do what I want. Why don't you just go and open your book or watch your illuminates and get the information yourself.

OpenStudy (anonymous):

OK STOP ARGUING this is dumb. We're not giving you the answers, sorry... These questions are all about lines, so I'll give you a few pointers. 1. All lines are of the form y=mx+b, where m is the slope and b is the initial value, aka y-intercept. 2. A horizontal line is of the form y=b, because the y-coordinate is always the same and there is no slope. 3. A vertical line is of the form x=b, because the x-coordinate is always the same and there is no slope. 4. If they say the line is parallel to another line, then they have the same slope. 5. If they say that 2 lines are perpendicular, then the slopes are negative reciprocals of each other. For example if one is 2/3, then the other is -3/2. If one is 4, then the other is -1/4.

OpenStudy (anonymous):

Hahahah, sad thing is I know the answers, and I can pass my math classes!

OpenStudy (anonymous):

Wow, you were not asking for the answers. Did you even read what you posted.

OpenStudy (anonymous):

Nova, please leave.

OpenStudy (anonymous):

Wow, I can also pass my English class with better vocabulary.

OpenStudy (anonymous):

Why not tell the person who is cursing to leave?

OpenStudy (anonymous):

You're both acting immaturely, no offense.

OpenStudy (anonymous):

Do you even realize that you can be expelled for cheating in online schools?

OpenStudy (anonymous):

I'm just trying to get some actual work done here.

OpenStudy (anonymous):

You don't get any work done here because she only wants the answers, not any help.

OpenStudy (anonymous):

You don't even know me..

OpenStudy (anonymous):

I know that you went back and deleted all of your postings where you were being rude and cursing to try to make yourself look better. I also know that instead of actually learning you are trying to take the easy way out and cheat. That is all I need to know about you.

OpenStudy (anonymous):

Obviously.